a monkey of mass 40 kg climbs on a rope which stands a maximum tension of 600n .in which case will the rope break?
 

Grade:11

2 Answers

Prabhat singh
5 years ago
Answer is that weight acting in downward direction will be weight=mg=40*9.8=392N whereas the maximum tension is 600N so there will be neither any case nor any situation in which the rope will break  
Swati malbhage
5 years ago
First we find the acceleration
F=mg+ma
F=40*10+40a
600=400+40a
40a=200
a=5m/s
Now,we put this acceleration in the equation,
S=ut+1/2at^2
10=1/2*5t^2
t^2=4
t=2 seconds